An investor deposited some money at 1.5% annual interest and two equal but larger amounts a time 2.2% and 2.4%. The total amount invested was $25000 and the total annual interest earned was $535. How much was invested at each rate?
Answer by addingup(3677)

Let the amount at 1.5% be x. Then:
0.015x+0.022((25000-x)/2)+0.024((25000-x)/2) = 535
0.022+0.024 = 0.046/2 = 0.023 Now we got rid of the two divisions. Rewrite the equation:
0.015x+0.023(25000-x) = 535
0.015x+575-0.023x = 535
-0.008x = -40 Divide both sides by -0.008 and remember -/- = +
x = 5000 This is how much was invested at 1.5% The rest, the problem says, was invested in equal amounts:
25000-5000 = 20000; 20000/2 = 10000 So, 10000 was invested each at 2.2% and 2.4%
Email me if you need more explanation.
:
Happy learning,
John